# FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models

### v3

@@ 4.2 Single- and Multi-Image Compressions
Experimental Setup. We evaluate our Factorized Features for image compression tasks against state-of-the-art methods. Following the Super-Resolution setup in Sec. 4.1, we train on the SA-1B dataset. To emphasize our representation, compression and decompression modules are initialized from pre-trained models [55], then fine-tuned end-to-end on 256×256 patches for 200k iterations using the AdamW optimizer [60] (learning rate=1e-5, batch size=16, betas=(0.9, 0.99)).
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 6.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
Rate-Distortion Performance Comparison. We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [56], [13], [98], [87], [19], [3], [44], [34], [66], [5], [71], and [32]. The classical image compression codec, VVC [80], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
In  Tab. 2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods. The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$. The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
Our experiments validate the effectiveness of FIPER in image compression, demonstrating substantial improvements across benchmarks (Kodak, CLIC, Tecnick). The multi-image compression strategy particularly excels with larger image sets, highlighting strong scalability. Additionally, our method achieves competitive latency, significantly improving compression efficiency while balancing quality and speed.
Table 4: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in Eq. 8.
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline [12] 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
4.3 Ablation Studies
Influence of SR Priors in Image Compression. We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness. TCM-HAT refers to using the original HAT [16] instead of our SR Module in Fig. 6b. TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly. Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Fields.
Effectiveness of Factorized Features Design. We conduct experiments to verify our multi-frequency modulation. The quantitative performance reported on single-image regression is shown in Tab. 4, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
@@ 5 Conclusion
Supplementary Material
Appendix A Overview
This supplementary material provides additional implementation details and experimental results to complement our main manuscript. Specifically, we first elaborate on the background of factor fields and learned image compression in Appendix B, which were omitted in the main paper due to space limitations. Next, in Appendix D, we present comprehensive details about the benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper. Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in Appendix H.
Appendix B Preliminaries
In this section, we briefly introduce the background of the factor fields [12], and Learned Image Compression along with their properties.
Factor Fields.
The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently [11, 27, 68, 26, 7, 9], while [12] first proposed a unified representation called factor fields. For a 1D signal $s(x)$, the factor fields formulation is:
$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (11)
where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation. The coefficient and basis are queried by sampling with the coordinates $x$. This generalizes to Q-dimensional signals:
$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (12)
where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
Learned Image Compression.
@@ Appendix B Preliminaries
$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},$
$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}),$
where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}})$. The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is then quantized to $\hat{z}$, where a factorized density model $\varphi$ is used to encode quantized $\hat{z}$ as $p_{\hat{z}\mid\varphi}(\hat{z}\mid\varphi)=\prod_{j}\left(p_{z_{j}\mid\varphi}(\varphi)*\mathcal{U}\left(-\frac{1}{2},\frac{1}{2}\right)\right)(\hat{z}_{j})$ where $j$ specifies the position or each element of each signal. The derivation of the equation can be found in [2]. Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
$L$ $=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x})$ (15)
@@
$+\lambda\cdot\mathcal{D}\left({x},\hat{{x}}\right)$
where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow [55], modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
Appendix C Differences from Factor Fields.
We explicitly highlight how sawtooth-based sampling and multi-frequency modulation extend the original factor fields. In our 1D and 2D toy examples as shown in Figs. 9 and 10, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$. Factor fields already improve accuracy over such baselines, but our introduction of $\psi$ and $\alpha$ further captures both high- and low-frequency details. Specifically, forcing $\alpha=4$ compels the network to learn broader frequency ranges and avoid noise from the fixed modulation. To confirm these advantages, we trained smaller SR models and compared them with factor fields in Tab. 5. The results illustrate that our multi-frequency constraint yields significant gains even in simpler configurations, underscoring its importance for high-frequency detail preservation and overall reconstruction quality.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Vanilla fields (b) Factor fields [10] (c) Ours with $\psi$ only (d) Ours with both $\psi$ and $\alpha$
Figure 9: 1D toy example.
@@
Table 5: Comparison with Factor Fields.
@@ Appendix D Implementation Details
The architecture of our SR networks is shown in Fig. 11. Firstly, for Coefficient Backbones from various pre-trained models like [15, 91, 17], we replace the output upsampling layer (mostly pixel shuffle) with convolution of channel $D=256$, i.e., for input image with resolution $64\times 64$, the original pre-trained models process hidden states in $64\times 64$, and then upsample the output to $256\times 256$ for final result, where we replace the upsampling with convolution and output the final hidden state $X_{coeff}\in R^{256\times 64\times 64}$.
@@
Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by Eq. 17:
@@
Tab. 6 shows quantitative comparison with state-of-the-art methods. All VSR experiments were conducted using bicubic 4X downsampling. The training dataset includes the REDS[69] and Vimeo-90K[89] datasets, while the testing dataset comprices REDS4[69], Vid4[53], and Vimeo-90K-T[89].
@@
As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [55], as shown in Fig. 12(more details can be bound in [55]. What we do here is that we use the output from the last main layer and convert it to match the hidden channel of the SR Module.
@@
Following VSR, we use transformer blocks to merge the bases. The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis. For the merged basis compression, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [55] with input and out channels change. After the Basis Merging Transformer is trained, we follow the same training setting for Basis compression while randomly sampling 1 to 24 images for a batch and setting only the Basis compression module trainable.
Appendix E Merged-Basis Strategy & SR Priors.
The SR prior accelerates detail recovery by leveraging high-resolution knowledge while basis merging exploits shared structures for efficiency. Since merging is introduced after, the compression model is pre-trained, and improvements stem not only from the SR prior but also from the well-initialized basis merging. When trained from scratch (see Tab. 7), the merging transformer tends to find a universal solution rather than image-specific structures. Additionally, merging inevitably causes information loss, prompting the coefficient branch to encode more image-specific details. This mutual offset between SR prior and basis merging ultimately strengthens our approach.
@@
Appendix F Encoding/Decoding Times.
Fig. 5(c) processes each of the $M$ images through the Coefficient Backbone, Basis Swin Transformer, and Basis Merging Transformer, then compresses and transmits the merged basis. In Tab. 2, different $M$ values primarily affect the run time of the Basis Merging Transformer. Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in Tab. 8. Notably, Tab. 2 assumes each image’s basis is compressed independently, which can be confusing. In practice, a merged basis is processed only once, so averaged encoding/decoding time is more realistic.
@@
Table 8: Comparison of Encoding/Decoding Times.
Appendix G Inference Time Comparison for SR.
We provide the inference time in Tab. 9, compared with extensive SR models. The inference time of different components of FIPER is also provided.
HAT-S HAT HAT-L ATD DAT RGT HAT-F-S (Ours) HAT-F(Ours) HAT-L-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
Inference Time (s) 0.0387 0.0398 0.0778 0.0592 0.0594 0.0612 0.0671 0.0716 0.1084 0.0400 0.0132 0.0184
Table 9: Inference time comparison for SR.

### v4

@@ 5.2 Single- and Multi-Image Compressions
Experimental Setup. Image-compression experiments follow the SR protocol (section˜5.1). Models are pretrained on ImageNet and fine-tuned for 200k steps on $\times256$ crops. Compression/decompression networks start from TCM weights [62] and are optimized end-to-end with AdamW (lr 1e-5, batch 16, $\beta$ = 0.9/0.99) [72]. Integrating our SR module into TCM yields TCM-HAT-F and TCM-HAT-L-F; TCM-HAT-F-multi is the multi-image variant.
Rate-Distortion Performance Comparison. We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [63], [13], [116], [101], [20], [3], [51], [40], [79], [5], [84], and [36]. The classical image compression codec, VVC [93], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution $\times1200$, and CLIC Professional Validation, is shown in fig.˜7. TCM-HAT-L-F achieves a -21.09% BD-Rate vs. VTM (Table 2), surpassing earlier work. Multi-image mode (TCM-HAT-F-multi) scales well, reaching -20.97% at M=24M{=}24: sending a shared basis curbs Coefficient-to-Basis error, though distortion grows slightly as basis capacity is stretched. Across Kodak, CLIC, and Tecnick, FIPER consistently boosts rate-distortion and maintains competitive latency.
@@ 6 Conclusion
Acknowledgements. This research was funded by the National Science and Technology Council, Taiwan, under Grants NSTC 112-2222-E-A49-004-MY2, 113-2628-E-A49-023-, and Academia Sinica under Grant AS- CDA-110-M09. The authors are grateful to Google, NVIDIA, and MediaTek Inc. for their generous donations, and we would also like to express our sincere gratitude to Jia-Wei Liao for valuable assistance in paper writing and revision. Yu-Lun Liu acknowledges the Yushan Young Fellow Program by the MOE in Taiwan.
Appendix A Overview
This supplementary material provides additional implementation details and experimental results to complement our main manuscript. Our FIPER framework explicity models multi-scale patterns and captures periodic features effectively, as visualized in fig.˜8.
@@ Appendix A Overview
In the following, we dive into the details. Specifically, we first elaborate on coordinate transformation, the background of learned image compression, and a comparison between single- and multi-image compression in appendix˜B, which were omitted in the main paper due to space limitations. Following this, we discuss the differences with factor fields [12] in appendix˜C. Next, in appendix˜D, we present comprehensive details about the experiments and benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper. Then, we provide ablation studies regarding different components of our pipeline, validating the design in appendix˜E. Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in appendix˜H.
Appendix B Method Details
In this section, we provide the details of Sawtooth coordinate transformation and downsampling inspired by it, the background of Learned Image Compression, and comparison between single- and multi-image compression.
B.1 Sawtooth Coordinate Transformation and Downsampling
Sawtooth transformation
Sawtooth transformation is formulated as $\gamma(x)=x\mod k,\ k\in R$.We can easily observe that such transformation implicitly captures patch-like frequency information as shown in fig.˜9a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
Downsampling based on Sawtooth transformation
To reconstruct a $256\times 256$ image, we formulate our Factorized Features as
which is the same as eq.˜7 in the main paper.
In practice, the bases are of shape $C\times H\times W$ with $N=6$. We set $H=W=\{256,\ 128,\ 64,\ 32,\ 16,\ 8\}$ and $C=24$ for all bases. Thus, we sample each basis with $\gamma_{i}(x)=x\mod k_{i}\ \text{where}\ x\in\{0,...,255\}^{2}$ and $k=\{256,128,64,32,16,8\}$. This formulation can be interpreted as we sample the same feature on different spatial location of a image, e.g., when using $b_{1}\in R^{24\times 128\times 128}$ with $k_{1}=128$, $x=\{0,0\}$ and $x=\{128,0\}$ will be extracting $b_{1}(0,0)$, which is also visualized in fig.˜9a. Therefore, we want to aggregate the information appearing in different location of the image, where these spatial locations will be sampled at the same location on the basis; that is, we want the information of $I(0,0)$ and $I(128,0)$ to be aggregated in $b_{1}(0,0)$ in previous example. This way, we can explicitly model repetitve structure in different scale, and such frequency modeling subsequently enhances the reconstruction.
Based on the analysis, as visualized in fig.˜9c, we integrate Sawtooth-Aware Downsample layer into Basis Transformer. Such operation is essentially the inverse of Sawtooth transformation, i.e., we merge the feature together with Sawtooth-Aware Downsample, decode the feature to basis, and then sample the basis to different spatial location via Sawtooth transformation.
B.2 Learned Image Compression.
@@ B.1 Sawtooth Coordinate Transformation and Downsampling
$\{F_{\text{mean}},F_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
$\hat{y}$ $=\{Q(y_{0}-\mu_{0})+\mu_{0},.,Q(y_{t}-\mu_{t})+\mu_{t}\},$
[…]
where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},F_{\text{mean}})$. The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is then quantized to $\hat{z}$, where a factorized density model $\varphi$ is used to encode quantized $\hat{z}$ as $p_{\hat{z}\mid\varphi}(\hat{z}\mid\varphi)=\prod_{j}\left(p_{z_{j}\mid\varphi}(\varphi)*\mathcal{U}\left(-\frac{1}{2},\frac{1}{2}\right)\right)(\hat{z}_{j})$ where $j$ specifies the position or each element of each signal. The derivation of the equation can be found in [2]. Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $F_{\text{mean}}$ and $F_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These $\mu_{t}$ are derived from earlier quantized slices and $F_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
[…]
@@
[…]
where R(y^)R($\hat${y}) and R(z^)R($\hat${z}) denote bit rates, D(x,x^)D($\hat{x}$) is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow [62], modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
B.3 Comparison between single- and multi-image compression
Figure 10: The illustration of our image-compression framework (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Factorized Features and Super-Resolution (SR) Module from section˜4.2 in the main paper as information-recovery prior. (c) Expanding on (b), we introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
We provide visual comparison of single- and multi-image compression in fig.˜10b and fig.˜10c, respectively, with traditional compression method in fig.˜10a. First, we have single image compression(SIC), the same as fig.˜5b in the main paper. The goal of the SIC pipeline is to validate two aspects: (1) that Factorized Features, which are designed to represent images, can be applied to different tasks; (2) that leveraging super‐resolution (SR) priors improves reconstruction quality. Note that we directly adopt the SR module’s architecture to also reduce overall complexity.
Next, we derive multi-image compression (MIC) as in section˜4.3.1 of the main paper. In traditional signal processing, a shared set of bases is often used to represent diverse sources. We adopt this property to further reduce compression redundancy. Specifically, our approach transmits (1) image features, which are used to decode coefficients, and (2) the bases themselves. By contrast, single image compression in fig.˜10b transmits only image features. Intuitively, transmitting the bases provides additional information that can improve reconstruction quality. However, directly transmitting all bases would significantly increase the bit rate. To address this, we merge the bases into a single set before transmission, thereby enhancing reconstruction while minimizing extra transmission costs.
Appendix C Differences from Factor Fields
We explicitly highlight how Sawtooth-based sampling and Multi-frequency Modulation extend the original factor fields. In 2D toy examples as shown in fig.˜11, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$. factor fields already improve accuracy over such baselines, but our introduction of $\psi$ and $\alpha$ further captures both high- and low-frequency details. Specifically, forcing $\alpha=4$ compels the network to learn broader frequency ranges and avoid noise from the fixed modulation. To confirm these advantages, we trained smaller SR models and compared them with factor fields in table˜3. The results illustrate that our multi-frequency constraint yields significant gains even in simpler configurations, underscoring its importance for high-frequency detail preservation and overall reconstruction quality.
@@ Appendix C Differences from Factor Fields
Table 3: Comparison with Factor Fields. HAT-FA-Small denotes replace our Factorized Features with factor fields.
@@ Appendix D Implementation Details
The architecture of our SR networks is shown in fig.˜12. Firstly, for Coefficient Backbones from various pre-trained models like [15, 109, 18], we replace the output upsampling layer (mostly pixel shuffle) with convolution of channel $D=256$, i.e., for input image with resolution $64\times 64$, the original pre-trained models process hidden states in $64\times 64$, and then upsample the output to $256\times 256$ for final result, where we replace the upsampling with convolution and output the final hidden state $X_{coeff}\in R^{256\times 64\times 64}$.
@@
Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by eq.˜15:
@@
table˜4 shows quantitative comparison with state-of-the-art methods. All VSR experiments were conducted using bicubic 4X downsampling. The training dataset includes the REDS[82] and Vimeo-90K[103] datasets, while the testing dataset comprices REDS4[82], Vid4[60], and Vimeo-90K-T[103].
@@
As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [62], as shown in fig.˜13(more details can be bound in [62]. What we do here is that we use the output from the last main layer and convert it to match the hidden channel of the SR Module.
@@
Following VSR, we use transformer blocks to merge the bases. The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis, i.e., if we have 16 bases of $W=H=64$ to merge into one, we rearrange the $64\times 64$ tokens as batch dimension, use the $16+1$ as sequence dim, and finally treat the additional 11(CLS) token as output. For the merged basis compression and transmission, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [62] with input and out channels change. After the Basis Merging Transformer is trained, we follow the same training setting for Basis compression while randomly sampling 1 to 24 images for a batch and setting only the Basis compression module trainable.
Appendix E Ablation Studies
E.1 Merged-Basis Strategy & SR Priors.
The SR prior accelerates detail recovery by leveraging high-resolution knowledge while basis merging exploits shared structures for efficiency. Since merging is introduced after, the compression model is pre-trained, and improvements stem not only from the SR prior but also from the well-initialized basis merging. When trained from scratch (see table˜5), the merging transformer tends to find a universal solution rather than image-specific structures. Additionally, merging inevitably causes information loss, prompting the coefficient branch to encode more image-specific details. This mutual offset between SR prior and basis merging ultimately strengthens our approach.
@@ Appendix E Ablation Studies
E.2 Encoding/Decoding Times.
Fig. 5(c) processes each of the $M$ images through the Coefficient Backbone, Basis Swin Transformer, and Basis Merging Transformer, then compresses and transmits the merged basis. In Tab. 2, different $M$ values primarily affect the run time of the Basis Merging Transformer. Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in table˜6. Notably, Tab. 2 assumes each image’s basis is compressed independently, which can be confusing. In practice, a merged basis is processed only once, so averaged encoding/decoding time is more realistic.
@@
BD-Rate $\%$ 27.96 2.70 -10.11 -16.61 -19.88 -20.97
Table 6: Comparison of Encoding/Decoding Times.
E.3 Inference Time Comparison for SR.
We provide the inference time in table˜7, compared with extensive SR models. The inference time of different components of FIPER is also provided.
ATD DAT RGT HAT-S HAT HAT-L HAT-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
Inference Time (s) 0.0592 0.0594 0.0612 0.0387 0.0398 0.0778 0.0716 0.0400 0.0132 0.0184
PSNR ont Set5 33.14 33.15 33.16 32.92 33.18 33.30 33.53 - - -
Table 7: Inference time comparison for SR.
E.4 Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness. TCM-HAT refers to using the original HAT [16] instead of our SR Module in fig.˜10b. TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly. Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Features.
Method Kodak CLIC Tecnick
@@
E.5 Effectiveness of Factorized Features Design.
We conduct experiments to verify our multi-frequency modulation. The quantitative performance reported on single-image regression is shown in table˜9, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
Table 9: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in eq.˜8. Note that here we test on single-image regression
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline [12] 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
E.6 Light-weight Model Comparison
To demonstrate the effectiveness of our method even under computation-constrained scenarios, we integrate our Factorized Features framework to state-of-the-art lightweight super-resolution models and observe consistent improvement as shown in table˜10.
Method Latency (ms) #FLOPs (G) #Params (K) Set5 PSNR / SSIM Set14 PSNR / SSIM B100 PSNR / SSIM Urban100 PSNR / SSIM Manga109 PSNR / SSIM
SwinIR-lt 222.9 63.6 930 32.44 / 0.8976 28.77 / 0.7858 27.69 / 0.7406 26.47 / 0.7980 30.92 / 0.9151
ELAN-lt 18.0 54.1 640 32.43 / 0.8975 28.78 / 0.7858 27.69 / 0.7406 26.54 / 0.7982 30.92 / 0.9150
OmniSR 22.5 50.9 792 32.49 / 0.8988 28.78 / 0.7859 27.71 / 0.7415 26.64 / 0.8018 31.02 / 0.9150
SRFormer-lt 287.2 62.8 873 32.51 / 0.8988 28.82 / 0.7872 27.73 / 0.7422 26.67 / 0.8032 31.17 / 0.9165
ATD-lt 189.7 100.1 769 32.63 / 0.8998 28.89 / 0.7886 27.79 / 0.7440 26.97 / 0.8107 31.48 / 0.9198
HiT-SRF 82.1 58.0 866 32.55 / 0.8999 28.87 / 0.7880 27.75 / 0.7432 26.80 / 0.8069 31.26 / 0.9171
ASID-D8 61.8 49.6 748 32.57 / 0.8990 28.89 / 0.7898 27.78 / 0.7449 26.89 / 0.8096
MambaIR-lt 55.8 84.6 924 32.42 / 0.8977 28.74 / 0.7847 27.68 / 0.7400 26.52 / 0.7983 30.94 / 0.9135
MambaIRV2-lt 153.4 75.6 790 32.51 / 0.8992 28.84 / 0.7878 27.75 / 0.7426 26.82 / 0.8079 31.24 / 0.9182
RDN 66.0 1309.2 22271 32.47 / 0.8990 28.81 / 0.7871 27.72 / 0.7419 26.61 / 0.8028 31.00 / 0.9151
RCAN 52.2 917.6 15592 32.63 / 0.9002 28.87 / 0.7889 27.77 / 0.7436 26.82 / 0.8087 31.22 / 0.9173
CATANet (CVPR’25) 102.4 49.3 535 32.58 / 0.8998 28.90 / 0.7880 27.75 / 0.7427 26.87 / 0.8081 31.31 / 0.9183
ESC (ICCV’25) 21.9 149.2 968 32.68 / 0.9011 28.93 / 0.7902 27.80 / 0.7447 27.07 / 0.8144 31.54 / 0.9207
CATANet-F (Ours) 110.3 56.7 941 32.69 / 0.9017 28.93 / 0.7906 27.82 / 0.7444 27.05 / 0.8133 31.62 / 0.9214
ESC-F (Ours) 24.6 73.4 952 32.74 / 0.9032 28.99 / 0.7931 27.86 / 0.7483 27.11 / 0.8158 31.78 / 0.9233
Table 10: Quantitative comparison on benchmark datasets for $\times 4$ lightweight image super-resolution.
E.7 Task-Specific Overfitting
To confirm that our Factorized Features are not overfitted to a single task, we further apply the same representation to other low-level vision problems such as motion deblurring in table˜11 and image dehazing in table˜12, without any task-specific modification. The consistent improvements across these tasks demonstrate that our features generalize well and capture transferable visual priors.
Metric / Method DeblurGAN-v2 SRN DMPHN SDWNet MPRNet MIMO-UNet+ DeepRFT+ MAXIM-3S Stripformer MSDI-net Restormer NAFNet FFTformer GRL-B MLWNet (CVPR’24) MLWNet-F (Ours)
PSNR 29.55 30.26 31.20 31.26 32.66 32.45 33.23 32.86 33.08 33.28 33.57 33.69 34.21 33.93 33.83 34.40
SSIM 0.934 0.934 0.945 0.966 0.959 0.957 0.963 0.961 0.962 0.964 0.966 0.967 0.969 0.968 0.968 0.971
Table 11: Performance comparison on the GoPro dataset for motion deblurring.
Metric / Method DehazeNet AOD-Net MSBDN FFA-Net AECR-Net DeHamer PMNet DehazeFormer TaylorFormer LH-Net MITNet PGH2Net (AAAI’24) ConvIR (TPAMI’24) ConvIR-F (Ours)
PSNR 13.84 13.14 15.37 14.39 15.80 16.62 16.79 16.29 16.66 18.87 16.97 17.02 17.45 18.12
SSIM 0.43 0.41 0.49 0.45 0.47 0.56 0.51 0.51 0.56 0.561 0.606 0.61 0.802 0.823
Table 12: Performance comparison on the Dense-Haze dataset for image dehazing.
E.8 Analysis of Model Hyperparameters
We further conduct a comprehensive study on three key hyperparameters of our model: the number of frequency components $N$, the frequency modulation $\alpha$, and the basis transformation $\gamma$ in table˜13 and table˜14. These factors collectively control the expressiveness and efficiency of the proposed Factorized Features.
NN (Number of Frequency Components) 1 1 2 6 6 6 6 10
$\alpha$ $\{1\}$ $\{1,4,16,64\}$ $\{1,4,16,64\}$ $\{1\}$ $\{1,4\}$ $\{1,4,16,64\}$ $\{1,4,16,64,256,1024\}$ $\{1,4,16,64\}$
Set5 PSNR 32.62 32.64 32.70 32.62 32.71 32.74 32.69 32.73
Set5 SSIM 0.9018 0.9020 0.9029 0.9020 0.9027 0.9032 0.9024 0.9033
Table 13: Ablation on the number of frequency components $N$ and scaling coefficients $\alpha$. We select $N=6$ for balanced quality and training efficiency, as $N=10$ yields marginal gains but converges slower.
Set5 Performance / $\gamma$ Sawtooth Sin Triangular
PSNR 32.71 32.69 32.71
SSIM 0.9027 0.9026 0.9023
Table 14: Comparison of alternative transformation bases $\gamma$.
These results show that our design choices provide a good trade-off between representation flexibility and efficiency, while maintaining stable performance across a wide range of hyperparameter configurations.
Appendix F Failure Cases
Failure cases typically involve extremely fine textures—details that are absent in the low-resolution (LR) input but appear in the super-resolved (SR) output. However, this has been a shared inherent limitation of super-resolution methods to date. For example, in top row of fig.˜6, some white structures are prevent in HR but completely absent in LR.
Appendix G Future Generalization
For image generation, for example, we can integrate our Factorized Features into the VAE decoder of a diffusion model. For video generation, we can employ our mergeable basis and leverage its properties to enhance structural coherence.